OpenStudy (anonymous):

Assuming an efficiency of 22.20%, calculate the actual yield of magnesium nitrate formed from 129.8 g of magnesium and excess copper(II) nitrate. Mg + Cu(NO3)2 -->Mg(NO3)2 + Cu

OpenStudy (aaronq):

Finding the theoretical yield: Use the stoichiometric coefficients to find moles produced. e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) so if you have moles of B= 2, how many moles of C did you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_c=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. then: Actual yield = % yield*theoretical yield